OpenStudy (anonymous):

Traveling at 45 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 240 miles from Orlando to Miami?

OpenStudy (anonymous):

How many miles do you travel in an hour?

OpenStudy (anonymous):

45

OpenStudy (kash_thesmartguy):

180 mins or 3 hours

OpenStudy (anonymous):

What do you mean?

OpenStudy (kash_thesmartguy):

It will take him 180 mins to go from Miami to Orlando

OpenStudy (anonymous):

No not correct

OpenStudy (anonymous):

The answer choices dont have 180 minutes

OpenStudy (anonymous):

I need help working the problem out.

OpenStudy (anonymous):

If you can travel 45 miles in one hour how many hours will it take you to go 240 miles?

OpenStudy (anonymous):

45 miles in one hour...90 miles in 2 hours....135 miles in 3 hours....etc....etc...

OpenStudy (anonymous):

5

OpenStudy (anonymous):

or.....divide 45 into 240 what do you get?

OpenStudy (anonymous):

45 into 240= 0.1875

OpenStudy (mathstudent55):

Here is the formula for speed: \(speed = \dfrac{distance}{time} \) we can abbreviate is as: \(s = \dfrac{d}{t} \) Your question is how many minutes it takes, so you are looking for time. We can solve our equation for t, time. \(s = \dfrac{d}{t} \) \(st = d\) \(t = \dfrac{d}{s} \) The time is the distance divided by the speed. Now plug in the known amounts your problem gives you. \(t = \dfrac{d}{s} \) \(t = \dfrac{240~miles}{45~mph} \) \(t = 5.3333... ~hours\) Now you need to convert from hours to minutes. An hour is 60 minutes. Since an hour is larger than a minute, in the same amount of time there are more minutes than hours. When you go from hours to minutes, the number will become larger. That means you multiply hours by 60 to get minutes. \(5.333... ~hours \times 60 ~\dfrac{minutes}{hour} = 320~minutes\)

OpenStudy (anonymous):

Thank you once again!

OpenStudy (mathstudent55):

You're welcome.